OpenStudy (anonymous):

What is the sum of the first 30 terms of the series made from the sequence in 0.3, 0.1, 0? Find the 40th term of the sequence 3, 6, 9,....?

OpenStudy (anonymous):

The 40th term of the sequence is 120...since it goes up by 3 each time. 40(3)=120

OpenStudy (anonymous):

ok..what about number 1?

OpenStudy (campbell_st):

looks like an arithmetic series with d = -0.2 a = 0.3 and n = 30 just substitute into the formula \[S_{30} = \frac{n}{2}[ 2a + (n -1)d]\] and evaluate

OpenStudy (campbell_st):

the 2nd question also appears to be arithmetic a = 3, d = 3 and n = 40 use the formula \[T_{40} = a + (n -1) d\]

OpenStudy (anonymous):

30/2 [2(0.3)+(30-1)-0.2 ? @campbell_st

OpenStudy (campbell_st):

not quite its \[S_{30} = \frac{30}{2}[2\times 0.3 + (30 - 1) \times (- 0.2)]\]